# SPEED: Scalable, Precise, and Efficient Concept Erasure for Diffusion Models

### v1

@@ Abstract
Erasing concepts from large-scale text-to-image (T2I) diffusion models has become increasingly crucial due to the growing concerns over copyright infringement, offensive content, and privacy violations. However, existing methods either require costly fine-tuning or degrade image quality for non-target concepts (i.e., prior) due to inherent optimization limitations. In this paper, we introduce SPEED, a model editing-based concept erasure approach that leverages null-space constraints for scalable, precise, and efficient erasure. Specifically, SPEED incorporates Influence-based Prior Filtering (IPF) to retain the most affected non-target concepts during erasing, Directed Prior Augmentation (DPA) to expand prior coverage while maintaining semantic consistency, and Invariant Equality Constraints (IEC) to regularize model editing by explicitly preserving key invariants during the T2I generation process. Extensive evaluations across multiple concept erasure tasks demonstrate that SPEED consistently outperforms existing methods in prior preservation while achieving efficient and high-fidelity concept erasure, successfully removing 100 concepts within just 5 seconds. Our code and models are available at: https://github.com/Ouxiang-Li/SPEED.
Figure 1: Three characteristics of our proposed method, SPEED. (a) Scalable: SPEED seamlessly scales from single-concept to large-scale multi-concept erasure (e.g., 100 celebrities) without additional design. (b) Precise: SPEED precisely removes the target concept (e.g., Snoopy) while preserving the semantic integrity for non-target concepts (e.g., Hello Kitty and SpongeBob). (c) Efficient: SPEED can immediately erase 100 concepts within 5 seconds, achieving a ×$\times 350$ speedup over the state-of-the-art (SOTA) method.
11footnotetext: Equal Contributions.22footnotetext: Corresponding author.
@@ 1 Introduction
Large-scale text-to-image (T2I) diffusion models [24, 56, 57, 39, 49, 23] have facilitated significant breakthroughs in generating highly realistic and contextually consistent images simply from textual descriptions [11, 46, 16, 7, 50, 44, 13]. Alongside these advancements, concerns have also been raised regarding copyright violations [10, 54, 30], offensive content [51, 65, 67], and privacy concerns [8, 64, 31]. To mitigate ethical and legal risks in generation, it is often necessary to prevent the model from generating certain concepts, a process termed concept erasure [29, 17, 66]. However, removing target concepts without carefully preserving the semantics of non-target concepts can introduce unintended artifacts, distortions, and degraded image quality  [17, 42, 51, 66], compromising the model’s reliability and usability. Therefore, beyond ensuring the effective removal of target concepts (i.e., erasure efficacy), concept erasure should also maintain the original semantics of non-target concepts (i.e., prior preservation).
@@
In this context, recent methods seek to balance between erasure efficacy and prior preservation, broadly categorized into two paradigms: training-based [29, 37, 35] and editing-based [18, 19]. The training-based paradigm fine-tunes T2I diffusion models to achieve concept erasure, incorporating an additional regularization term in training for prior preservation. In contrast, the editing-based paradigm is training-free. It jointly optimizes erasure and preservation objectives for target and non-target concepts in a closed-form solution, obtaining direct modifications over model parameters (e.g., projection weights in cross-attention layers [49]). This efficiency also allows editing-based methods to seamlessly extend to multi-concept erasure without additional designs.
@@
However, as the number of target concepts increases, current editing-based methods struggle to balance between erasure efficacy and prior preservation. It can be attributed to the growing conflicts among erasure and preservation objectives, making such trade-offs increasingly difficult. Moreover, these methods rely on weighted least squares optimization [18, 19], inherently imposing a non-zero lower bound on preservation error (Appx. B.2). In multi-concept settings, this accumulation of preservation errors distorts non-target knowledge, further degrading generative quality.
To address the above limitations, we propose Scalable, Precise, and Efficient Concept Erasure for Diffusion Models (SPEED) (see Fig. 1), an editing-based method incorporating null-space constraints. Here, we define “null-space” as the subspace that does not alter the feature representations of non-target concepts. By projecting the model parameter updates for erasing concepts onto the null space of non-target concepts, SPEED can minimize the preservation error close to zero without compromising erasure efficacy. A key challenge for SPEED lies in constructing an effective null space based on a set of non-target concepts (i.e., retain set). We observe that the existing baseline with null-space constraints [14] confronts a fundamental dilemma during concept erasure: While a narrow retain set limits the prior coverage, increasing the diversity of the retain set restricts the determination of the null space due to rank saturation, thereby compromising prior preservation (see Fig. 2). In this light, we introduce Prior Knowledge Refinement, a suite of techniques that alleviates optimization difficulty and broadens prior coverage.
@@
Particularly, to mitigate optimization difficulty, we propose Influence-based Prior Filtering (IPF), which selectively retains non-target concepts from the retain set that are most affected by the erasure process. These concepts serve as anchors for defining the null space, providing meaningful constraints for preservation while maintaining sufficient flexibility for model editing. To further prevent excessive prior knowledge reduction, we introduce Directed Prior Augmentation (DPA) to enhance prior coverage, which expands the selected retain set after IPF with augmented samples. Instead of simply introducing random noise, we generate semantically consistent variations of preserved concepts, mitigating excessive preservation over meaningless concepts. Furthermore, we incorporate Invariants Equality Constraint (IEC), which identifies critical invariants in the generation process and enforces equality regularization to explicitly preserve them. We evaluate SPEED on three representative concept erasure tasks, i.e., few-concept, multi-concept, and implicit concept erasure, where it consistently exhibits superior prior preservation across all erasure tasks. Overall, our contributions are summarized as follows:
@@
  • We introduce Prior Knowledge Refinement to construct a well-defined null space that remains meaningful and tractable for editing. By integrating Influence-based Prior Filtering, Directed Prior Augmentation, and Invariants Equality Constraint, it balances rank saturation and prior coverage effectively with precise concept erasure.
@@
Figure 2: Impact of the increasing number of non-target concepts in retain set. The baseline [14] maintains prior preservation with a small retain set (e.g., Refer to caption), but as the retain set grows, rank saturation occurs, leading to gradually degraded semantics in non-target generations (e.g., Refer to caption and Refer to caption). In contrast, our method achieves consistent prior preservation by optimizing the retain set.
@@ 2 Related Works
Concept erasure. Current T2I diffusion models inevitably involve unauthorized and NSFW (Not Safe For Work) generations due to the noisy training data from web [52, 53]. Apart from applying additional filters or safety checkers [47, 41, 48], prevailing methods modify diffusion model parameters to erase specific target concepts, mainly categorized into two paradigms. The training-based paradigm fine-tunes model parameters with specific erasure objectives [29, 17, 66] and additional regularization terms [29, 37, 35]. In contrast, the editing-based paradigm edits model parameters using a closed-form solution to facilitate efficiency in concept erasure. For example, UCE [18] modifies model weights by balancing both erasure and preservation error through a weighted least squares objective and RECE [19] iteratively derives new target concept embeddings. These methods can erase numerous concepts within seconds, demonstrating superior efficiency in practice.
Null-space constraints. The null space of a matrix, a fundamental concept in linear algebra, refers to the set of all vectors that the matrix maps to the zero vector. The null-space constraints are first applied to continual learning (CL) by projecting gradients onto the null space of uncentered covariances from previous tasks [60]. Subsequent studies [36, 61, 63, 28, 32] further explore and extend the application of null space in CL. In model editing, AlphaEdit [14] restricts model weight updates onto the null space of preserved knowledge, effectively mitigating trade-offs between editing and preservation. The null-space constraints also apply to various other tasks, e.g., machine unlearning [9] and MRI reconstruction [15], offering great promise for editing-based concept erasure.
@@ 3.1 Concept Erasure in Closed-Form Solution
In T2I diffusion models, each concept is encoded by a set of text tokens via CLIP [45], which are then aggregated into a single concept embedding $\boldsymbol{c}\in\mathbb{R}^{d_{0}}$. For concept erasure, there are two sets of concepts: the erasure set $\mathbf{E}$ and the retain set $\mathbf{R}$. The erasure set consists of $N_{E}$ target concepts to be removed, denoted as $\mathbf{E}=\{\boldsymbol{c}_{1}^{(i)}\}_{i=1}^{N_{E}}$. The retain set includes $N_{R}$ non-target concepts that should be preserved during editing, denoted as $\mathbf{R}=\{\boldsymbol{c}_{0}^{(i)}\}_{i=1}^{N_{R}}$. To effectively erase each target concept $\boldsymbol{c}_{1}^{(i)}\in\mathbf{E}$ (e.g., Snoopy), it is mapped onto an anchor concept $\boldsymbol{c}_{*}^{(i)}$ that retains general semantics while avoiding the original concept-specific information (e.g., Dog). The paired erasure set is denoted as: $\mathbf{\tilde{E}}=\{(\boldsymbol{c}_{1}^{(i)},\boldsymbol{c}_{*}^{(i)})\}_{i=1}^{N_{E}}$.
@@
For editing-based methods [42, 18, 19], concept embeddings from the erasure set $\mathbf{\tilde{E}}$ and the retain set $\mathbf{R}$ are first organized into three structured matrices: $\mathbf{C}_{1},\mathbf{C}_{*}\in\mathbb{R}^{d_{0}\times N_{E}}$ and $\mathbf{C}_{0}\in\mathbb{R}^{d_{0}\times N_{R}}$, which represents the stacked embeddings of the target, anchor, and non-target concepts, respectively. To derive a closed-form solution for concept erasure, existing methods typically optimize a perturbation $\boldsymbol{\Delta}$ to model parameters $\mathbf{W}$, balancing between erasure efficacy and prior preservation. For example, UCE [18] formulates concept erasure as a weighted least squares problem:
$\boldsymbol{\Delta}_{\text{UCE}}=\mathop{\text{arg\ min}}\limits_{\boldsymbol{\Delta}}\underbrace{\|(\mathbf{W}+\boldsymbol{\Delta})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}}_{\text{$e_{1}$}}+\underbrace{\|\boldsymbol{\Delta}\mathbf{C}_{0}\|^{2}}_{\text{$e_{0}$}},$ (1)
@@
where the erasure error $e_{1}$ ensures that each target concept is mapped onto its corresponding anchor concept and the preservation error $e_{0}$ minimizes the impact on non-target priors. This formulation provides a closed-form solution to obtain $\boldsymbol{\Delta}_{\text{UCE}}$ (see Appx. B.1) for parameter updates, achieving computationally efficient optimization. However, as the number of target concepts increases, the accumulated preservation errors $e_{0}$ across multiple erased concepts would amplify the distortion on non-target knowledge, further degrading prior preservation.
3.2 Null-Space Constraints
To mitigate the limitation of weighted optimization in prior preservation, SPEED integrates null-space constraints to achieve model editing [14]. These constraints restrict the solution space of $\boldsymbol{\Delta}$ within the null space of $\mathbf{C}_{0}$ to ensure that parameter updates do not interfere with non-target concepts, enforcing $e_{0}$ close to zero without a tradeoff between $e_{0}$ and $e_{1}$. To obtain $\boldsymbol{\Delta}$, we first perform singular value decomposition (SVD) on $\mathbf{C}_{0}\mathbf{C}_{0}^{\top}$***Considering $\mathbf{C}_{0}\in\mathbb{R}^{d_{0}\times N_{R}}$ may have high dimensionality depending on $N_{R}$, we compute the null space by performing SVD on $\mathbf{C}_{0}\mathbf{C}_{0}^{\top}\in\mathbb{R}^{d_{0}\times d_{0}}$, as both $\mathbf{C}_{0}$ and $\mathbf{C}_{0}\mathbf{C}_{0}^{\top}$ share the same null space. and construct a projection matrix $\mathbf{P}$:
$\left\{\mathbf{U},\mathbf{\Lambda},\mathbf{U}^{\top}\right\}=\text{SVD}\left(\mathbf{C}_{0}\mathbf{C}_{0}^{\top}\right),\quad\mathbf{P}=\hat{\mathbf{U}}\hat{\mathbf{U}}^{\top},$ (2)
where $\hat{\mathbf{U}}$ is the matrix composed of eigenvectors w.r.t. the zero singular values in $\mathbf{\Lambda}$. This projection matrix $\mathbf{P}$ can map $\boldsymbol{\Delta}$ onto the null space of $\mathbf{C}_{0}$, i.e., $\|\left(\boldsymbol{\Delta}\mathbf{P}\right)\mathbf{C}_{0}\|^{2}=0$. Thus, the objective becomes:
$\left(\boldsymbol{\Delta}\mathbf{P}\right)_{\text{Null}}=$ $\mathop{\text{arg\ min}}\limits_{\boldsymbol{\Delta}}\underbrace{\|(\mathbf{W}+\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}}_{\text{$e_{1}$}}$

### v2

@@ Abstract
Erasing concepts from large-scale text-to-image (T2I) diffusion models has become increasingly crucial due to the growing concerns over copyright infringement, offensive content, and privacy violations. In scalable applications, fine-tuning-based methods are time-consuming to precisely erase multiple target concepts, while real-time editing-based methods often degrade the generation quality of non-target concepts due to conflicting optimization objectives. To address this dilemma, we introduce SPEED, an efficient concept erasure approach that directly edits model parameters. SPEED searches for a null space, a model editing space where parameter updates do not affect non-target concepts, to achieve scalable and precise erasure. To facilitate accurate null space optimization, we incorporate three complementary strategies: Influence-based Prior Filtering (IPF) to selectively retain the most affected non-target concepts, Directed Prior Augmentation (DPA) to enrich the filtered retain set with semantically consistent variations, and Invariant Equality Constraints (IEC) to preserve key invariants during the T2I generation process. Extensive evaluations across multiple concept erasure tasks demonstrate that SPEED consistently outperforms existing methods in non-target preservation while achieving efficient and high-fidelity concept erasure, successfully erasing 100 concepts within only 5 seconds. Our code and models are available at: https://github.com/Ouxiang-Li/SPEED.
Figure 1: Three characteristics of our proposed concept erasure method for diffusion models, SPEED. (a) Scalable: SPEED seamlessly scales from single-concept to large-scale multi-concept erasure (e.g., 100 celebrities) without additional design. (b) Precise: SPEED precisely removes the target concept (e.g., Snoopy) while preserving the semantics for non-target concepts (e.g., Hello Kitty and SpongeBob). (c) Efficient: SPEED immediately erases 100 concepts within 5 seconds, achieving new state-of-the-art (SOTA) performance with a $350\times$ speedup over competitive methods.
@@ 1 Introduction
Text-to-image (T2I) diffusion models Ho et al. (2020); Song et al. (2020a; b); Nichol & Dhariwal (2021); Rombach et al. (2022); Ho & Salimans (2022) have facilitated significant breakthroughs in generating highly realistic and contextually consistent images simply from textual descriptions Dhariwal & Nichol (2021); Ramesh et al. (2021); Gal et al. (2022); Betker et al. (2023); Ruiz et al. (2023); Podell et al. (2023); Esser et al. (2024). Alongside these advancements, concerns have also been raised regarding copyright violations Cui et al. (2023); Shan et al. (2023), offensive content Schramowski et al. (2023); Yang et al. (2024b); Zhang et al. (2025), and privacy concerns Carlini et al. (2023); Yang et al. (2023). To mitigate ethical and legal risks in generation, it is often necessary to prevent the model from generating certain concepts, a process termed concept erasure Kumari et al. (2023); Gandikota et al. (2023); Zhang et al. (2024a). However, removing target concepts without carefully preserving the semantics of non-target concepts can introduce unintended artifacts, distortions, and degraded image quality Gandikota et al. (2023); Orgad et al. (2023); Schramowski et al. (2023); Zhang et al. (2024a), compromising the model’s usability. Therefore, beyond ensuring the effective removal of target concepts (i.e., erasure efficacy), concept erasure should also maintain the original semantics of non-target concepts (i.e., prior preservation).
@@
In this context, recent methods strive to seek a balance between erasure efficacy and prior preservation, broadly categorized into two paradigms: training-based Kumari et al. (2023); Lyu et al. (2024); Lu et al. (2024a) and editing-based Gandikota et al. (2024); Gong et al. (2025). The training-based paradigm fine-tunes diffusion models to achieve concept erasure, incorporating an additional regularization into the training objective for prior preservation. In contrast, the editing-based paradigm avoids additional fine-tuning by directly modifying model parameters (e.g., projection weights in cross-attention layers Rombach et al. (2022)), with such modifications derived from a closed-form objective that jointly accounts for erasure and preservation. This efficiency also facilitates editing-based methods to extend to multi-concept erasure without additional designs seamlessly.
@@
However, as the number of target concepts increases, current editing-based methods Gandikota et al. (2024); Gong et al. (2025) struggle to balance between erasure efficacy and prior preservation. This can be attributed to the growing conflicts between erasure and preservation objectives, making such trade-offs increasingly difficult. Moreover, these methods rely on weighted least squares optimization, inherently imposing a non-zero lower bound on preservation error (see Appx. B.2). In multi-concept settings, this accumulation of preservation errors gradually distorts non-target knowledge, thereby degrading prior preservation. To address the above limitations, we propose Scalable, Precise, and Efficient Concept Erasure for Diffusion Models (SPEED) (see Fig. 1), an editing-based method incorporating null-space constraints. Specifically, we search for the null space of prior knowledge, a model editing space where parameter updates do not affect the feature representations of non-target concepts. By projecting the model parameter updates for concept erasure onto such null space, SPEED can minimize the preservation error to zero without compromising erasure efficacy, thereby enabling scalable and precise concept erasure without affecting non-target concepts.
Figure 2: Semantic degradation with increasing non-target concepts in retain set. The baseline null-space constrained method Fang et al. (2024) preserves non-target semantics given a small retain set Refer to caption. However, as the retain set expands, the corresponding matrix approaches higher rank, making null space estimation increasingly inaccurate (see Eq. 4) with inevitable approximation errors, thereby degrading prior semantics Refer to caption Refer to caption.
The key contribution of SPEED lies in defining an effective null space from a set of non-target concepts (i.e., retain set). We observe that the existing baseline with null-space constraints Fang et al. (2024) confronts a fundamental dilemma during concept erasure: While a small retain set limits the coverage of prior knowledge, enlarging the retain set makes it increasingly difficult to identify an accurate null space. This difficulty arises because a large retain set causes the corresponding feature matrix to approach full rank, necessitating the estimation of its null space to ensure sufficient degrees of freedom for optimization (i.e., concept erasure). However, this estimation inevitably introduces semantic degradation to the retain set and deteriorates prior preservation in Fig. 2.
@@
In this light, we introduce Prior Knowledge Refinement, a suite of techniques that strategically and selectively refine the retain set to mitigate the semantic degradation in searching for the null space. Particularly, we propose Influence-based Prior Filtering (IPF), which first quantifies the influence of concept erasure on each non-target concept. It then prunes the retain set by removing minimally affected concepts, preventing the correlation matrix from approaching full rank and thus maintaining an accurate null space. Subsequently, to further enhance prior preservation over the resulting retain set, we propose Directed Prior Augmentation (DPA), which expands the retain set with directed, semantically consistent perturbations to improve retain coverage. In addition, we incorporate Invariant Equality Constraints (IEC) to preserve specific representations, such as the [SOT] token, that should remain unchanged during editing. IEC enforces equality constraints on such invariants to regularize the retaining of essential generation properties. We evaluate SPEED on three representative concept erasure tasks, i.e., few-concept, multi-concept, and implicit concept erasure, where it consistently exhibits superior prior preservation across all erasure tasks. Overall, our contributions can be summarized as follows:
@@
  • We introduce Prior Knowledge Refinement to construct an accurate null space over the retain set for effective editing. Leveraging three complementary techniques, IPF, DPA, and IEC, our method balances semantic degradation and retain coverage, enabling precise and scalable concept erasure.
@@ 2 Related Works
Concept erasure. Current T2I diffusion models inevitably involve unauthorized and offensive generations due to the noisy training data from web Schuhmann et al. (2021; 2022). Apart from applying additional filters or safety checkers Rando et al. (2022); Betker et al. (2023); Rao (2023), prevailing methods modify diffusion model parameters to erase specific target concepts, mainly categorized into two paradigms. The training-based paradigm fine-tunes model parameters with specific erasure objectives Kumari et al. (2023); Gandikota et al. (2023); Zhang et al. (2024a); Zhao et al. (2024b); Huang et al. (2024); Kim et al. (2024); Zhang et al. (2024b); Zhao et al. (2024a) and additional regularization terms Kumari et al. (2023); Lyu et al. (2024); Lu et al. (2024a). In contrast, the editing-based paradigm edits model parameters using a closed-form solution to facilitate efficiency in concept erasure Orgad et al. (2023); Gandikota et al. (2024); Gong et al. (2025). These methods can erase numerous concepts within seconds, demonstrating superior efficiency in practice. Beyond parameter modification, non-parametric methods (e.g., external modules and sampling interventions) have also been explored Schramowski et al. (2023); Wang et al. (2024b); Yoon et al. (2024); Jain et al. (2024); Lee et al. (2025b; a), but they are fragile in open-source settings.
Null-space constraints. The null space of a matrix, a fundamental concept in linear algebra, refers to the set of all vectors that the matrix maps to the zero vector. The null-space constraints are first applied to continual learning by projecting gradients onto the null space of uncentered covariances from previous tasks Wang et al. (2021). Subsequent studies Lu et al. (2024b); Wang et al. (2024a); Yang et al. (2024a); Kong et al. (2022); Lin et al. (2022) further explore and extend the application of null space in continual learning. In model editing, AlphaEdit Fang et al. (2024) restricts model weight updates onto the null space of preserved knowledge, effectively mitigating trade-offs between editing and preservation. Null-space constraints also apply to various tasks, e.g., machine unlearning Chen et al. (2024), MRI reconstruction Feng et al. (2023), and image restoration Wang et al. (2022), offering promise for editing-based concept erasure.
@@ 3 Problem Formulation
In T2I diffusion models, each concept is encoded by a set of text tokens via CLIP Radford et al. (2021), which are then aggregated into a single concept embedding $\bm{c}\in\mathbb{R}^{d_{0}}$. For concept erasure, there are two sets of concepts: the erasure set $\mathbf{E}$ and the retain set $\mathbf{R}$. The erasure set consists of $N_{E}$ target concepts to be removed, denoted as $\mathbf{E}=\{\bm{c}_{1}^{(i)}\}_{i=1}^{N_{E}}$. The retain set includes $N_{R}$ non-target concepts that should be preserved during editing, denoted as $\mathbf{R}=\{\bm{c}_{0}^{(j)}\}_{j=1}^{N_{R}}$. To enable efficient erasure efficacy for $\mathbf{E}$ and prior preservation for $\mathbf{R}$, we first formulate a closed-form editing objective in Sec. 3.1, and enhance it with null-space constrained optimization in Sec. 3.2.
@@ 3.1 Concept Erasure in Closed-Form Solution
To effectively erase each target concept $\bm{c}_{1}^{(i)}\in\mathbf{E}$ (e.g., Snoopy), it is specified to be mapped onto an anchor concept $\bm{c}_{*}^{(i)}$ that shares general semantics (e.g., Dog), termed as an anchor set $\mathbf{A}=\{\bm{c}_{*}^{(i)}\}_{i=1}^{N_{E}}$. For editing-based methods Orgad et al. (2023); Gandikota et al. (2024); Gong et al. (2025), concept embeddings from the erasure set $\mathbf{E}$, anchor set $\mathbf{A}$, and retain set $\mathbf{R}$ are first organized into three structured matrices: $\mathbf{C}_{1},\mathbf{C}_{*}\in\mathbb{R}^{d_{0}\times N_{E}}$ and $\mathbf{C}_{0}\in\mathbb{R}^{d_{0}\times N_{R}}$, representing the stacked embeddings of target, anchor, and non-target concepts, respectively. To derive a closed-form solution for concept erasure, existing methods typically optimize a perturbation $\bm{\Delta}$ to model parameters $\mathbf{W}$, balancing between erasure efficacy and prior preservation. For example, UCE Gandikota et al. (2024) formulates concept erasure as a weighted least squares problem:
$\bm{\Delta}_{\text{UCE}}=\mathop{\text{arg\ min}}\limits_{\bm{\Delta}}\underbrace{\|(\mathbf{W}+\bm{\Delta})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}}_{\text{$e_{1}$}}+\underbrace{\|\bm{\Delta}\mathbf{C}_{0}\|^{2}}_{\text{$e_{0}$}},$ (1)
@@
where the erasure error $e_{1}$ ensures that each target concept is mapped onto its corresponding anchor concept and the preservation error $e_{0}$ minimizes the impact on non-target concepts, and $\|\cdot\|^{2}$ denotes the sum of the squared elements in the matrix (i.e., Frobenius norm). This formulation provides a closed-form solution $\bm{\Delta}_{\text{UCE}}$ (see Appx. B.1) for parameter updates, achieving computationally efficient optimization. However, as the number of target concepts increases, the accumulated preservation errors $e_{0}$, which prove to share a non-zero bound from Appx. B.2, across multiple target concepts would amplify the distortion on non-target knowledge and degrade prior preservation.
3.2 Apply Null-Space Constraints
To address the limitation of weighted optimization in prior preservation, SPEED incorporates null-space constraints Wang et al. (2021); Fang et al. (2024) to achieve prior-preserved model editing by forcing $e_{0}=0$. The null space of $\mathbf{C}_{0}$ consists of all vectors $\bm{v}$ such that $\bm{v}\mathbf{C}_{0}=\mathbf{0}$. Restricting the parameter update $\bm{\Delta}$ to this space ensures that such updates do not interfere with non-target concepts.